OpenStudy (anonymous):

Determine the largest multiple of 6 which divides 14! Explain your work.

jimthompson5910 (jim_thompson5910):

14! = 14*13*12*11*10*9*8*7*6*5*4*3*2*1 Rewrite the right side in the form of 6k where k is some number

jimthompson5910 (jim_thompson5910):

make sure k is as large as possible

OpenStudy (anonymous):

im having trouble understanding. are you saying 12 is the answer?

jimthompson5910 (jim_thompson5910):

no, 12 isn't the answer

jimthompson5910 (jim_thompson5910):

can you rearrange 14*13*12*11*10*9*8*7*6*5*4*3*2*1 into the form of 6k

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

is there a 6 anywhere in that list?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

ok pull that out and rewrite into the form 6k where k is some number

OpenStudy (anonymous):

ok i pull out 6k, then what?

OpenStudy (anonymous):

im not getting this

jimthompson5910 (jim_thompson5910):

14! = 14*13*12*11*10*9*8*7*6*5*4*3*2*1 14! = 6*14*13*12*11*10*9*8*7*5*4*3*2*1 14! = 6*(14*13*12*11*10*9*8*7*5*4*3*2*1)

OpenStudy (anonymous):

and then i just calculate it?

jimthompson5910 (jim_thompson5910):

yeah calculate everything after the 6*

OpenStudy (anonymous):

is the answer within the parenthesis or is it the answer after i multiply that all by six

jimthompson5910 (jim_thompson5910):

the result after you multiply all that out in the parenthesis

jimthompson5910 (jim_thompson5910):

because that will give you the multiple of 6

jimthompson5910 (jim_thompson5910):

that's a factor of 14!

OpenStudy (anonymous):

yea but is the answer what's in the parenthesis? or that times 6

jimthompson5910 (jim_thompson5910):

whatever is in the parenthesis, multiply that out

OpenStudy (anonymous):

and that's the answer whatever k is

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np